# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 6.1.2 Results and Analysis
TIRG18’CVPR ResNet18 LSTM $14.87$ $34.66$ $18.26$ $37.89$ $19.08$ $39.62$ $17.40$ $37.39$
VAL20’CVPR ResNet50 LSTM $22.53$ $44.00$ $22.38$ $44.15$ $27.53$ $51.68$ $24.15$ $46.61$
MAAF20’Arxiv ResNet50 LSTM $23.8$ $48.6$ $21.3$ $44.2$ $27.9$ $53.6$ $24.3$ $48.8$
TRACE21’AAAI ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.7$ $61.23$ $29.08$ $54.7$
RTIC21’Arxiv ResNet50 LSTM $19.4$ $43.51$ $16.93$ $38.36$ $21.58$ $47.88$ $19.3$ $43.25$
DCNet21’AAAI ResNet50 Glove $28.95$ $56.07$ $23.95$ $47.30$ $30.44$ $58.29$ $27.78$ $53.89$
CoSMo21’CVPR ResNet50 LSTM $25.64$ $50.3$ $24.90$ $49.18$ $29.21$ $57.46$ $26.58$ $52.31$
HFCA21’ACMMM ResNet50 LSTM $26.2$ $51.2$ $22.4$ $46.0$ $29.7$ $56.4$ $26.1$ $51.2$
CLVC-Net21’SIGIR ResNet50 LSTM $29.85$ $56.47$ $28.75$ $54.76$ $33.50$ $64.00$ $30.70$ $58.41$
CIRPLANT21’ICCV ResNet152 LSTM $14.38$ $34.66$ $13.64$ $33.56$ $16.44$ $38.34$ $14.82$ $35.52$
CIRPLANT W/ OSCAR21’ICCV ResNet152 LSTM $17.45$ $40.41$ $17.53$ $38.81$ $21.64$ $45.38$ $18.87$ $41.53$
JPM(VAL+MSE)21’ACMMM ResNet18 LSTM $21.27$ $43.12$ $21.88$ $43.3$ $25.81$ $50.27$ $22.98$ $45.59$
JPM(VAL+Tri)21’ACMMM ResNet18 LSTM $21.38$ $45.15$ $22.81$ $45.18$ $27.78$ $51.70$ $23.99$ $47.34$
FashionVLP22’CVPR ResNet50 BERT $32.42$ $60.29$ $31.89$ $58.44$ $38.51$ $68.79$ $34.27$ $62.51$
SAC22’WACV ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.70$ $61.23$ $29.08$ $54.70$
ERR22’TIP ResNet50 LSTM $30.02$ $55.44$ $25.32$ $49.87$ $33.20$ $60.34$ $29.51$ $55.22$
CRR22’ACMMM ResNet101 BiGRU $30.41$ $57.11$ $30.73$ $58.02$ $33.67$ $64.48$ $31.60$ $59.87$
Progressive22’SIGIR ResNet50 BERT $33.22$ $59.99$ $46.17$ $68.79$ $46.46$ $73.84$ $41.98$ $67.54$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $38.18$ $64.50$ $48.63$ $71.54$ $52.32$ $76.90$ $46.37$ $70.98$
MCR22’TMM ResNet50 LSTM $26.20$ $51.20$ $22.40$ $46.00$ $29.70$ $56.40$ $26.10$ $51.20$
NEUCORE23’NIPS-W ResNet BiGRU $27.00$ $53.79$ $22.84$ $45.00$ $29.63$ $56.65$ $26.45$ $51.75$
ARTEMIS22’ICLR ResNet50 LSTM $27.34$ $51.71$ $21.05$ $49.87$ $27.34$ $44.18$ $24.43$ $48.59$
ARTEMIS22’ICLR ResNet50 BiGRU $27.16$ $52.40$ $21.78$ $43.64$ $29.20$ $54.83$ $26.05$ $50.29$
AMC23’TOMM ResNet50 LSTM $31.73$ $59.25$ $36.21$ $66.60$ $30.67$ $59.08$ $32.87$ $61.64$
ComqueryFormer23’TMM Swin-T BERT $33.86$ $61.08$ $35.57$ $62.19$ $42.07$ $69.30$ $37.17$ $64.19$
RankUn23’Arxiv CLIP(RN50) CLIP $34.8$ $60.22$ $45.21$ $69.06$ $47.68$ $74.85$ $42.5$ $68.04$
CRN23’TIP Swin-T(base) LSTM $30.24$ $57.61$ $29.83$ $55.54$ $33.91$ $64.04$ $31.36$ $59.06$
CRN23’TIP Swin-T(Large) LSTM $32.67$ $59.30$ $30.27$ $56.97$ $37.74$ $65.94$ $33.56$ $60.74$
SSN23’Arxiv CLIP(ViT-B/32) CLIP $34.36$ $60.78$ $38.13$ $61.83$ $44.26$ $69.025$ $38.92$ $63.89$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $45.22$ $69.66$ $52.60$ $72.52$ $56.14$ $77.10$ $51.32$ $73.09$
MCEMiner24’TIP ResNet50 LSTM $33.23$ $59.16$ $26.15$ $50.87$ $33.83$ $61.40$ $31.07$ $57.14$
CaLa-CLIP4Cir24’SIGIR $32.93$ $56.82$ $39.20$ $60.13$ $39.16$ $63.83$ $37.10$ $60.26$
CaLa-ARTEMIS24’SIGIR $40.13$ $66.88$ $46.86$ $67.28$ $49.87$ $74.11$ $45.62$ $69.42$
CaLa-BLIP2cir24’SIGIR $42.38$ $66.08$ $46.76$ $68.16$ $50.93$ $73.42$ $46.69$ $69.22$
CSS-Net24’KBS ResNet50 RoBERTa $33.65$ $63.16$ $35.96$ $61.96$ $42.65$ $70.70$ $37.42$ $65.27$
SPIRIT24’TOMM CLIP(RN50) CLIP $43.83$ $68.86$ $52.50$ $74.19$ $56.60$ $79.25$ $50.98$ $74.10$
LIMN24’TPAMI ResNet50 LSTM $35.6$ $62.37$ $34.69$ $59.81$ $40.64$ $68.33$ $36.98$ $63.50$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $57.63$ $78.56$ $62.14$ $80.38$ $66.15$ $85.73$ $61.97$ $81.56$
Uncertainty24’ICLR ResNet50 RoBERTa $30.60$ $57.46$ $31.54$ $58.29$ $37.37$ $68.41$ $33.17$ $61.39$
@@
SDQUR24’TCSVT BLIP2 $56.87$ $76.50$ $49.93$ $73.33$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM ResNet50 LSTM $30.19$ $58.80$ $29.39$ $55.69$ $37.66$ $64.97$ $32.26$ $59.76$
SPRC24’ICLR BLIP2(ViT-L) $49.18$ $72.43$ $55.64$ $73.89$ $59.35$ $78.58$ $54.92$ $74.97$
@@
Progressive(stage1)22’SIGIR ResNet50 BERT $5.75$ $13.04$ $11.73$ $21.98$ $11.22$ $21.93$ $9.57$ $18.98$
TABLE III: Comparison of Different Methods on Fashion200K & Shoes Datasets
@@
TIRG18’CVPR ResNet18 LSTM $14.1$ $42.5$ $63.8$ $40.1$
VAL20’CVPR MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
LBF(small)20’CVPR Faster-RCNN $16.26$ $46.9$ $71.73$ $44.96$
LBF(big)20’CVPR Faster-RCNN $17.8$ $48.35$ $68.5$ $44.83$
VAL20’CVPR MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
JSVM20’ECCV MobileNet LSTM $19.0$ $52.1$ $70.0$ $47.0$
TRACE21’AAAI ResNet50 GRU $18.5$ $51.73$ $77.28$
MAAF20’Arxiv ResNet50 LSTM $18.94$
GSCMR21’TIP Faster-RCNN BiGRU $21.57$ $52.84$ $70.12$ $48.18$
MAN21’ICASSP ResNet18 BERT $17.1$ $47.9$ $68.1$ $44.37$
MAN21’ICASSP MobileNet BERT $22.3$ $54.5$ $74.1$ $50.3$
RTIC21’Arxiv ResNet50 LSTM $43.66$ $72.11$
ComposeAE21’WACV ResNet18 BERT $22.8$ $55.3$ $73.4$ $50.5$
DCNet21’AAAI ResNet50 Glove $46.89$ $67.56$ $53.82$ $79.33$
CoSMo21’CVPR ResNet18&ResNet50 LSTM $23.3$ $50.4$ $69.3$ $47.7$ $16.72$ $48.36$ $75.64$ $46.91$
HFCA21’ACMMM ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$ $17.85$ $50.95$ $77.24$ $48.68$
CLVC-Net21’SIGIR ResNet50 LSTM $22.6$ $53.0$ $72.2$ $49.3$ $17.64$ $54.39$ $79.47$ $50.50$
HCL21’MMAsia ResNet18 LSTM $23.48$ $54.03$ $73.71$ $50.41$
JPM(TIRG+MSE)21’ACMMM ResNet18 LSTM $19.8$ $46.5$ $66.6$ $44.3$
JPM(TIRG+Tri)21’ACMMM ResNet18 LSTM $17.7$ $44.7$ $64.5$ $42.3$
FashionVLP22’CVPR ResNet18&ResNet50 BERT $49.9$ $70.5$ $60.2$ $49.08$ $77.32$ $63.20$ $63.20$
SAC22’WACV ResNet50 GRU $18.5$ $51.73$ $77.28$ $49.17$
ERR22’TIP ResNet50 LSTM $50.88$ $70.60$ $19.87$ $55.96$ $79.58$ $51.80$
TriArea22’Sci.Rep. ResNet18+TF LSTM $17.7$ $46.8$ $66.2$ $43.6$
CRR22’ACMMM ResNet101 GRU&BiGRU $24.85$ $56.41$ $73.56$ $51.61$ $19.41$ $56.38$ $79.92$ $51.90$
MultiColSAP22’TMM ResNet50 LSTM $51.06$ $70.13$ $60.60$ $19.54$ $65.39$ $79.47$ $54.80$
MCR22’TMM ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$
Progressive22’SIGIR ResNet50 BERT $19.53$ $55.65$ $80.58$ $51.92$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $22.88$ $58.53$ $84.16$ $55.29$
ComqueryFormer23’TMM Swin-T BERT $52.2$ $72.2$ $62.2$
NEUCORE23’NIPS-W ResNet BiGRU $19.76$ $55.48$ $80.75$
ARTEMIS22’ICLR ResNet50 LSTM $17.60$ $51.05$ $76.85$ $48.50$
ARTEMIS22’ICLR ResNet50 BiGRU $18.72$ $53.11$ $79.31$ $50.38$
AMC23’TOMM ResNet50 LSTM $19.99$ $56.89$ $79.27$
CRN23’TIP Swin-T(base) LSTM $17.32$ $54.15$ $79.34$ $50.27$ $53.3$ $73.3$ $63.3$ $63.3$
CRN23’TIP Swin-T(Large) LSTM $18.92$ $54.55$ $80.04$ $51.17$ $53.5$ $74.5$ $64.0$ $64.0$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $58.05$
MCEMiner24’TIP ResNet18&ResNet50 LSTM $26.82$ $56.76$ $76.91$ $53.50$ $19.10$ $55.37$ $79.57$ $51.35$
CSS-Net24’KBS RoBERTa RoBERTa $22.2$ $50.5$ $69.7$ $47.5$ $20.13$ $56.81$ $81.32$ $52.75$
SPIRIT24’TOMM CLIP(RN50) CLIP $55.2$ $73.6$ $56.90$ $81.49$ $69.19$ $69.19$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $36.8$ $67.9$ $87.8$ $64.1$ $31.47$ $69.19$ $88.52$ $63.06$
LIMN24’TPAMI ResNet50 LSTM $57.30$ $82.7$
Uncertainty24’ICLR ResNet50 RoBERTa $21.8$ $52.1$ $70.2$ $48.0$ $18.41$ $53.63$ $79.84$ $50.63$
SDQUR24’TCSVT BLIP2 $30.14$ $68.30$ $88.30$ $62.25$
AlRet24’TMM ResNet18&ResNet50 LSTM $24.42$ $53.93$ $73.25$ $53.53$ $18.13$ $53.98$ $78.81$ $50.31$
AlRet24’TMM CLIP LSTM $21.02$ $55.72$ $80.77$ $52.50$
NSFSE24’TMM ResNet50 BiGRU $25.3$ $53.8$ $73.5$ $50.87$
NSFSE24’TMM ResNet152 BiGRU $24.9$ $54.3$ $73.4$ $50.87$
MANME24’TCSVT ResNet50 BiGRU $23.0$ $57.9$ $75.3$ $52.0$ $20.73$ $55.96$ $80.98$ $52.56$
@@
TIRG18’CVPR ResNet18 LSTM $73.70$ $46.60$
LBF(small)20’CVPR Faster-RCNN $67.26$ $50.31$
LBF(big)20’CVPR Faster-RCNN $79.20$ $55.69$
MAAF20’Arxiv ResNet50 LSTM $87.80$
GSCMR21’TIP Faster-RCNN BiGRU $81.81$ $58.74$
MAN21’ICASSP ResNet18 BERT $79.60$
MAN21’ICASSP MobileNet BERT $80.40$
HCL21’MMAsia ResNet18 LSTM $81.59$ $58.65$
JPM(TIRG+MSE)21’ACMMM ResNet18 LSTM $83.80$
JPM(TIRG+Tri)21’ACMMM ResNet18 LSTM $83.20$
CRR22’ACMMM ResNet101 BiGRU $85.84$

### v2

@@ 6.1.2 Results and Analysis
TIRG18’CVPR [20] ResNet18 LSTM $14.87$ $34.66$ $18.26$ $37.89$ $19.08$ $39.62$ $17.40$ $37.39$
VAL20’CVPR [48] ResNet50 LSTM $22.53$ $44.00$ $22.38$ $44.15$ $27.53$ $51.68$ $24.15$ $46.61$
MAAF20’Arxiv [49] ResNet50 LSTM $23.8$ $48.6$ $21.3$ $44.2$ $27.9$ $53.6$ $24.3$ $48.8$
TRACE21’AAAI [50] ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.7$ $61.23$ $29.08$ $54.7$
RTIC21’Arxiv [58] ResNet50 LSTM $19.4$ $43.51$ $16.93$ $38.36$ $21.58$ $47.88$ $19.3$ $43.25$
DCNet21’AAAI [128] ResNet50 Glove $28.95$ $56.07$ $23.95$ $47.30$ $30.44$ $58.29$ $27.78$ $53.89$
CoSMo21’CVPR [37] ResNet50 LSTM $25.64$ $50.3$ $24.90$ $49.18$ $29.21$ $57.46$ $26.58$ $52.31$
HFCA21’ACMMM [42] ResNet50 LSTM $26.2$ $51.2$ $22.4$ $46.0$ $29.7$ $56.4$ $26.1$ $51.2$
CLVC-Net21’SIGIR [55] ResNet50 LSTM $29.85$ $56.47$ $28.75$ $54.76$ $33.50$ $64.00$ $30.70$ $58.41$
CIRPLANT21’ICCV [21] ResNet152 LSTM $14.38$ $34.66$ $13.64$ $33.56$ $16.44$ $38.34$ $14.82$ $35.52$
CIRPLANT(OSCAR)21’ICCV [21] ResNet152 LSTM $17.45$ $40.41$ $17.53$ $38.81$ $21.64$ $45.38$ $18.87$ $41.53$
JPM(VAL+MSE)21’ACMMM [68] ResNet18 LSTM $21.27$ $43.12$ $21.88$ $43.3$ $25.81$ $50.27$ $22.98$ $45.59$
JPM(VAL+Tri)21’ACMMM [68] ResNet18 LSTM $21.38$ $45.15$ $22.81$ $45.18$ $27.78$ $51.70$ $23.99$ $47.34$
FashionVLP22’CVPR [56] ResNet50 BERT $32.42$ $60.29$ $31.89$ $58.44$ $38.51$ $68.79$ $34.27$ $62.51$
SAC22’WACV [45] ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.70$ $61.23$ $29.08$ $54.70$
ERR22’TIP [44] ResNet50 LSTM $30.02$ $55.44$ $25.32$ $49.87$ $33.20$ $60.34$ $29.51$ $55.22$
CRR22’ACMMM [59] ResNet101 BiGRU $30.41$ $57.11$ $30.73$ $58.02$ $33.67$ $64.48$ $31.60$ $59.87$
Progressive22’SIGIR [237] ResNet50 BERT $33.22$ $59.99$ $46.17$ $68.79$ $46.46$ $73.84$ $41.98$ $67.54$
Progressive22’SIGIR [237] CLIP(ViT-B/32) BERT $38.18$ $64.50$ $48.63$ $71.54$ $52.32$ $76.90$ $46.37$ $70.98$
MCR22’TMM [43] ResNet50 LSTM $26.20$ $51.20$ $22.40$ $46.00$ $29.70$ $56.40$ $26.10$ $51.20$
NEUCORE23’NIPS-W [46] ResNet BiGRU $27.00$ $53.79$ $22.84$ $45.00$ $29.63$ $56.65$ $26.45$ $51.75$
ARTEMIS22’ICLR [129] ResNet50 LSTM $27.34$ $51.71$ $21.05$ $49.87$ $27.34$ $44.18$ $24.43$ $48.59$
ARTEMIS22’ICLR [129] ResNet50 BiGRU $27.16$ $52.40$ $21.78$ $43.64$ $29.20$ $54.83$ $26.05$ $50.29$
AMC23’TOMM [66] ResNet50 LSTM $31.73$ $59.25$ $36.21$ $66.60$ $30.67$ $59.08$ $32.87$ $61.64$
ComqueryFormer23’TMM [238] Swin-T BERT $33.86$ $61.08$ $35.57$ $62.19$ $42.07$ $69.30$ $37.17$ $64.19$
RankUn23’Arxiv [100] CLIP(RN50) CLIP $34.8$ $60.22$ $45.21$ $69.06$ $47.68$ $74.85$ $42.5$ $68.04$
CRN23’TIP [51] Swin-T(base) LSTM $30.24$ $57.61$ $29.83$ $55.54$ $33.91$ $64.04$ $31.36$ $59.06$
CRN23’TIP [51] Swin-T(Large) LSTM $32.67$ $59.30$ $30.27$ $56.97$ $37.74$ $65.94$ $33.56$ $60.74$
SSN23’Arxiv [67] [67] CLIP(ViT-B/32) CLIP $34.36$ $60.78$ $38.13$ $61.83$ $44.26$ $69.025$ $38.92$ $63.89$
TG-CIR23’ACMMM [72] CLIP(ViT-B/16) CLIP $45.22$ $69.66$ $52.60$ $72.52$ $56.14$ $77.10$ $51.32$ $73.09$
MCEMiner24’TIP [71] ResNet50 LSTM $33.23$ $59.16$ $26.15$ $50.87$ $33.83$ $61.40$ $31.07$ $57.14$
CaLa-CLIP4Cir24’SIGIR [60] $32.93$ $56.82$ $39.20$ $60.13$ $39.16$ $63.83$ $37.10$ $60.26$
CaLa-ARTEMIS24’SIGIR [60] $40.13$ $66.88$ $46.86$ $67.28$ $49.87$ $74.11$ $45.62$ $69.42$
CaLa-BLIP2cir24’SIGIR [60] $42.38$ $66.08$ $46.76$ $68.16$ $50.93$ $73.42$ $46.69$ $69.22$
CSS-Net24’KBS [99] ResNet50 RoBERTa $33.65$ $63.16$ $35.96$ $61.96$ $42.65$ $70.70$ $37.42$ $65.27$
SPIRIT24’TOMM [61] CLIP(RN50) CLIP $43.83$ $68.86$ $52.50$ $74.19$ $56.60$ $79.25$ $50.98$ $74.10$
LIMN24’TPAMI [57] ResNet50 LSTM $35.6$ $62.37$ $34.69$ $59.81$ $40.64$ $68.33$ $36.98$ $63.50$
DQU-CIR24’SIGIR [130] CLIP(ViT-H/14) $57.63$ $78.56$ $62.14$ $80.38$ $66.15$ $85.73$ $61.97$ $81.56$
Uncertainty24’ICLR [101] ResNet50 RoBERTa $30.60$ $57.46$ $31.54$ $58.29$ $37.37$ $68.41$ $33.17$ $61.39$
@@
SDQUR24’TCSVT [65] BLIP2 $56.87$ $76.50$ $49.93$ $73.33$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM [39] ResNet50 LSTM $30.19$ $58.80$ $29.39$ $55.69$ $37.66$ $64.97$ $32.26$ $59.76$
SPRC24’ICLR [63] BLIP2(ViT-L) $49.18$ $72.43$ $55.64$ $73.89$ $59.35$ $78.58$ $54.92$ $74.97$
@@
Progressive(stage1)22’SIGIR [237] ResNet50 BERT $5.75$ $13.04$ $11.73$ $21.98$ $11.22$ $21.93$ $9.57$ $18.98$
TABLE III: Comparison of Different Methods on Fashion200K & Shoes Datasets.
@@
TIRG18’CVPR [20] ResNet18 LSTM $14.1$ $42.5$ $63.8$ $40.1$
VAL20’CVPR [48] MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
LBF(small)20’CVPR [40] Faster-RCNN $16.26$ $46.9$ $71.73$ $44.96$
LBF(big)20’CVPR [40] Faster-RCNN $17.8$ $48.35$ $68.5$ $44.83$
VAL20’CVPR [48] MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
JSVM20’ECCV [69] MobileNet LSTM $19.0$ $52.1$ $70.0$ $47.0$
TRACE21’AAAI [50] ResNet50 GRU $18.5$ $51.73$ $77.28$
MAAF20’Arxiv [49] ResNet50 LSTM $18.94$
GSCMR21’TIP[62] Faster-RCNN BiGRU $21.57$ $52.84$ $70.12$ $48.18$
MAN21’ICASSP [136] ResNet18 BERT $17.1$ $47.9$ $68.1$ $44.37$
MAN21’ICASSP [136] MobileNet BERT $22.3$ $54.5$ $74.1$ $50.3$
RTIC21’Arxiv [58] ResNet50 LSTM $43.66$ $72.11$
ComposeAE21’WACV [38] ResNet18 BERT $22.8$ $55.3$ $73.4$ $50.5$
DCNet21’AAAI [128] ResNet50 Glove $46.89$ $67.56$ $53.82$ $79.33$
CoSMo21’CVPR [37] ResNet18&ResNet50 LSTM $23.3$ $50.4$ $69.3$ $47.7$ $16.72$ $48.36$ $75.64$ $46.91$
HFCA21’ACMMM [42] ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$ $17.85$ $50.95$ $77.24$ $48.68$
CLVC-Net21’SIGIR [55] ResNet50 LSTM $22.6$ $53.0$ $72.2$ $49.3$ $17.64$ $54.39$ $79.47$ $50.50$
HCL21’MMAsia [52] ResNet18 LSTM $23.48$ $54.03$ $73.71$ $50.41$
JPM(TIRG+MSE)21’ACMMM [68] ResNet18 LSTM $19.8$ $46.5$ $66.6$ $44.3$
JPM(TIRG+Tri)21’ACMMM [68] ResNet18 LSTM $17.7$ $44.7$ $64.5$ $42.3$
FashionVLP22’CVPR [56] ResNet18&ResNet50 BERT $49.9$ $70.5$ $60.2$ $49.08$ $77.32$ $63.20$ $63.20$
SAC22’WACV [45] ResNet50 GRU $18.5$ $51.73$ $77.28$ $49.17$
ERR22’TIP [44] ResNet50 LSTM $50.88$ $70.60$ $19.87$ $55.96$ $79.58$ $51.80$
TriArea22’Sci.Rep. [44] ResNet18+TF LSTM $17.7$ $46.8$ $66.2$ $43.6$
CRR22’ACMMM [59] ResNet101 GRU&BiGRU $24.85$ $56.41$ $73.56$ $51.61$ $19.41$ $56.38$ $79.92$ $51.90$
MultiColSAP22’TMM [124] ResNet50 LSTM $51.06$ $70.13$ $60.60$ $19.54$ $65.39$ $79.47$ $54.80$
MCR22’TMM [43] ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$
Progressive22’SIGIR [237] ResNet50 BERT $19.53$ $55.65$ $80.58$ $51.92$
Progressive22’SIGIR [237] CLIP(ViT-B/32) BERT $22.88$ $58.53$ $84.16$ $55.29$
ComqueryFormer23’TMM [238] Swin-T BERT $52.2$ $72.2$ $62.2$
NEUCORE23’NIPS-W [46] ResNet BiGRU $19.76$ $55.48$ $80.75$
ARTEMIS22’ICLR [129] ResNet50 LSTM $17.60$ $51.05$ $76.85$ $48.50$
ARTEMIS22’ICLR [129] ResNet50 BiGRU $18.72$ $53.11$ $79.31$ $50.38$
AMC23’TOMM [66] ResNet50 LSTM $19.99$ $56.89$ $79.27$
CRN23’TIP [51] Swin-T(base) LSTM $17.32$ $54.15$ $79.34$ $50.27$ $53.3$ $73.3$ $63.3$ $63.3$
CRN23’TIP [51] Swin-T(Large) LSTM $18.92$ $54.55$ $80.04$ $51.17$ $53.5$ $74.5$ $64.0$ $64.0$
TG-CIR23’ACMMM [72] CLIP(ViT-B/16) CLIP $58.05$
MCEMiner24’TIP [71] ResNet18&ResNet50 LSTM $26.82$ $56.76$ $76.91$ $53.50$ $19.10$ $55.37$ $79.57$ $51.35$
CSS-Net24’KBS [99] RoBERTa RoBERTa $22.2$ $50.5$ $69.7$ $47.5$ $20.13$ $56.81$ $81.32$ $52.75$
SPIRIT24’TOMM [61] CLIP(RN50) CLIP $55.2$ $73.6$ $56.90$ $81.49$ $69.19$ $69.19$
DQU-CIR24’SIGIR [130] CLIP(ViT-H/14) $36.8$ $67.9$ $87.8$ $64.1$ $31.47$ $69.19$ $88.52$ $63.06$
LIMN24’TPAMI [57] ResNet50 LSTM $57.30$ $82.7$
Uncertainty24’ICLR [101] ResNet50 RoBERTa $21.8$ $52.1$ $70.2$ $48.0$ $18.41$ $53.63$ $79.84$ $50.63$
SDQUR24’TCSVT [65] BLIP2 $30.14$ $68.30$ $88.30$ $62.25$
AlRet24’TMM [39] ResNet18&ResNet50 LSTM $24.42$ $53.93$ $73.25$ $53.53$ $18.13$ $53.98$ $78.81$ $50.31$
AlRet24’TMM [39] CLIP LSTM $21.02$ $55.72$ $80.77$ $52.50$
NSFSE24’TMM [47] ResNet50 BiGRU $25.3$ $53.8$ $73.5$ $50.87$
NSFSE24’TMM [47] ResNet152 BiGRU $24.9$ $54.3$ $73.4$ $50.87$
MANME24’TCSVT [53] ResNet50 BiGRU $23.0$ $57.9$ $75.3$ $52.0$ $20.73$ $55.96$ $80.98$ $52.56$
@@
TIRG18’CVPR [20] ResNet18 LSTM $73.70$ $46.60$
LBF(small)20’CVPR [40] Faster-RCNN $67.26$ $50.31$
LBF(big)20’CVPR [40] Faster-RCNN $79.20$ $55.69$
MAAF20’Arxiv [49] ResNet50 LSTM $87.80$
GSCMR21’TIP [62] Faster-RCNN BiGRU $81.81$ $58.74$
MAN21’ICASSP [136] ResNet18 BERT $79.60$
MAN21’ICASSP [136] MobileNet BERT $80.40$
HCL21’MMAsia [52] ResNet18 LSTM $81.59$ $58.65$
JPM(TIRG+MSE)21’MM [68] ResNet18 LSTM $83.80$
JPM(TIRG+Tri)21’MM [68] ResNet18 LSTM $83.20$
CRR22’ACMMM [59] ResNet101 BiGRU $85.84$